OpenStudy (anonymous):

Challenge question: Give a proof that Pi is irrational. (NO need to prove transcendental if not necessary)

OpenStudy (anonymous):

proof by contradiction :)

OpenStudy (lgbasallote):

proof by bandwagon

Parth (parthkohli):

Let's assume that \(\pi\) is rational.\[\pi = {p \over q} \quad(\gcd{(p,q) = 1)} \]\[\pi q = p \]and now contradict.

Parth (parthkohli):

Wait... I think I am going wrong.

OpenStudy (anonymous):

I have to excuse myself- got to leave. MEDALS AND MENTIONS AND fan will be given TO anybody and everybody who CONTRIBUTES to the CORRECT answer

Parth (parthkohli):

A number is irrational if: 1) It has infinite number of decimal numbers after the decimal. 2) It shows no pattern after the decimal(there are no repeating digits) Pi possesses all the qualities shown above.

Parth (parthkohli):

I don't think I am good at proofs. :(

OpenStudy (experimentx):

the proof of irrationality of pi is not simple as other numbers http://www.proofwiki.org/wiki/Pi_is_Irrational

OpenStudy (anonymous):

\[\tan^{-1}x=x-\frac{x^3}{3!}.........\] put x=1 we get \[\frac{\pi}{4}=1-\frac{1}{3!}+\frac{1}{5!}-............................\]

OpenStudy (zzr0ck3r):

wow. thats pretty nutty.

OpenStudy (experimentx):

a friend of mine tried to formulate it using two concentric circles ... almost spent a week ... in the end he found that he was wrong.

OpenStudy (zzr0ck3r):

was a good week though

OpenStudy (experimentx):

crazy guy!!

OpenStudy (anonymous):

Now we must proof that the series converges and we will have it

OpenStudy (zzr0ck3r):

lol what?

OpenStudy (experimentx):

nowdays there's a concept called rational trigonometry ... http://en.wikipedia.org/wiki/Rational_trigonometry ... quite different from it. though i don't understand it ... might be what you are looking for

OpenStudy (anonymous):

This whole post is a waste of time. You know why? Because Pi is wrong: http://www.youtube.com/watch?v=jG7vhMMXagQ&feature=edu&list=PL5F03A9D6D278C5D9

OpenStudy (experimentx):

Pi is not really wrong ... if it would have been replaced by tau then things would have been much simpler. well ... to me doesn't make much difference.

OpenStudy (anonymous):

I think it can be proved with Taylor's series error, I saw a proof of e being irrational in Calculus I course with taylor's series.

OpenStudy (zzr0ck3r):

wow tau owns.

OpenStudy (anonymous):

man i have not that patience to read proofs about pi irrationality or e but sin 1 is better

OpenStudy (anonymous):

eem of course e's proof is much simpler than pi

OpenStudy (experimentx):

i forgot for sin(1) though ...

OpenStudy (mayankdevnani):

IT is proved that if x is non-zero and rational then this expression must be irrational. Since tan(π⁄4) = 1, it follows that π⁄4 is irrational and therefore that π is irrational.....

OpenStudy (anonymous):

@mayankdevnai This is mcuh more elegant - only if the statement " if x is non-zero and rational then this expression must be irrational" can be proven though.

OpenStudy (anonymous):

@mayankdevnani This is mcuh more elegant - only if the statement " if x is non-zero and rational then this expression must be irrational" can be proven though.

OpenStudy (mayankdevnani):

@Mikael oh!! i tried it by step by step...